OpenStudy (erniellyd28):

Let y=f(x) describe the temperature of a room that is kept at 73 degrees f for x hours. a) represent f symbolically and graphically over a 24 hour period for 0

OpenStudy (michele_laino):

please can you make a screenshot of the original text of the exercise?

OpenStudy (erniellyd28):

im not high tech pro so im not sure how to do screen shots

OpenStudy (erniellyd28):

im very sorry

OpenStudy (michele_laino):

I think that we have a constant temperature, so we get the subsequent ordewred pairs: \[\begin{gathered} \left( {0,73} \right) \hfill \\ \left( {4,73} \right) \hfill \\ \left( {8,73} \right) \hfill \\ \left( {12,73} \right) \hfill \\ \left( {24,73} \right) \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

oops.. I meant ordered* pairs

OpenStudy (michele_laino):

 I think that it is a possible table, wherein \(T\) stands for temperature please complete such table